A researcher wants to study the relationship between salary and gender. She randomly selects 368 individuals and determines their salary and gender. Can the researcher conclude that salary and gender are dependent?

Income Male Female Total
Below $25,000 57 60 117
$⁢25,000-$50,000 48 44 92
$⁢50,000-$75,000 48 45 93
Above $75,000 3131 35 66
Total 184 184 368

Copy Data

Step 1 of 8:

State the null and alternative hypothesis.

Step 2 of 8:

Find the expected value for the number of men with an income below $25,000. Round your answer to one decimal place.

Step 3 of 8:

Find the expected value for the number of women with an income of $⁢25,000-$50,000. Round your answer to one decimal place.

Step 4 of 8:

Find the value of the test statistic. Round your answer to three decimal places.

Step 5 of 8:

Find the degrees of freedom associated with the test statistic for this problem.

Step 6 of 8:

Find the critical value of the test at the 0.005 level of significance. Round your answer to three decimal places.

Step 7 of 8:

Make the decision to reject or fail to reject the null hypothesis at the 0.005 level of significance.

Step 8 of 8:

State the conclusion of the hypothesis test at the 0.005 level of significance.
Answer #1

1) The null and alternative hypothesis is

H0: There is no relationship between salary and gender.

H1: There is a relationship between salary and gender.

2)

The formula for expected value.

E = ( Row total*Column total) / Grand total

E1 = (184*117)/368 = 58.5

3)

E4 = (184*92)/368 = 46

4)

Test statistic is

\chi ^{2}=\sum \frac{(O-E)^{2}}{E}

O: Observed frequency
E: Expected frequency.
E = ( Row total*Column total) / Grand total

O E (O-E) (O-E)^2 (O-E)^2/E
57 58.5 -1.5 2.25 0.038462
60 58.5 1.5 2.25 0.038462
48 46 2 4 0.086957
44 46 -2 4 0.086957
48 46.5 1.5 2.25 0.048387
45 46.5 -1.5 2.25 0.048387
31 33 -2 4 0.121212
35 33 2 4 0.121212
Total 0.590

(o- E)2

5)

Degrees of freedom = ( Number of rows - 1 ) * ( Number of column - 1) = ( 4 - 1) * (2 - 1) = 3 * 1 = 3

6)

Level of significance = 0.005

Critical value = 12.838

( From chi-square table)

7)

Test statistic < critical value we fail to reject null hypothesis.

8)

Conclusion:

There is no relationship between salary and gender.